OpenStudy (anonymous):

s=r*theta (2.5=r*pi/3)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

What do you want done with these equations? Solve for a certain variable or graph it or what?

OpenStudy (anonymous):

i can't find r or at least i'm not quite sure how

OpenStudy (anonymous):

So you're saying that s=2.5, theta=pi/3, and you want to find r? If so, then:\[2.5=r*\frac{\pi}{3}\]Divide both sides by pi/3, which is same as multiplying by 3/pi (its reciprocal).\[2.5*\frac{3}{\pi}=r\]\[\frac{7.5}{\pi}=r\]

OpenStudy (anonymous):

Thanks So Much! That really helped, to be honest you did a better job at explaining than my pre cal teacher.

OpenStudy (anonymous):

No problem, I love helping people, especially when their teachers don't bother to.